f is diff at 1. find f '(1)


if lim h→0 f(1+h) / h =5.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear dinesh

lim h→0 f(1+h) / h =5

limit exist onlt when f(1+h) =0 at h=0

so f(1)=0

now

lim h→0 {f(1+h) -f(1)}/ {(1+h)-1)       =5

 which is   f'(1) =5


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin